The decomposition of N2O5 follows first-order kinetics based on the reaction below. 

 

2N2O5(g)        O2(g) + 4NO2(g) 

 

The rate constant of the reaction at 25ºC is 3.46 x 10-5 /s. How long (in seconds) would it take for an initial concentration of 0.050 M to decrease to 0.015 M? What is the half-life of the reaction? 


1
Expert's answer
2021-05-12T06:39:58-0400

The reaction is

"2N_2O_5\Rightarrow 4NO_2+O_2"


In question we have given that

Initial concentration"(C_0)" "=0.050" M

Final concentration"(C_t)" "=0.015" M

Rate constant(K) "=" "3.46\times10^{-5}" per second


Now for the first order reaction:


"T=\dfrac{2.303}{K}Log\dfrac{C_0}{C_t}"


"T=\dfrac{2.303}{3.46\times10^{-5}}Log\dfrac{0.050}{0.015}=3.46\times10^4" seconds


Now half life period for the first order reaction"=\dfrac{0.693}{K}"


"=\dfrac{0.693}{3.46\times10^{-5}}=2.00\times10^{4}" second